# The evolution of inharmonicity and noisiness in contemporary popular music

### v1

@@ 3.2.2 Peak prominence and noisiness
Features have been proposed to measure the ‘noisiness’ of the signal, such as spectral flatness (Peeters,, 2004) and AudioSpectrumFlatnessType (ISO,, 2001). However, the two features are sensitive to the overall envelope of the spectrum: music and pink noise result in comparable spectral flatness values, which defeats the purpose of measuring ‘noisiness’. We introduce a new metric derived from spectral flatness and AudioSpectrumFlatnessType that is robust to the overall spectral profile. We refer to this feature as peak prominence and to the inverse of peak prominence as noisiness. Section 7 details the elaboration of the metric.
@@ 3.2.3 Inharmonicity
Inharmonicity has been defined as ‘the departure in frequency from the harmonic modes of vibration’ (Young,, 1952), and ‘the deviation of a set of frequencies from an exact harmonic series’ (Campbell,, 2001). A further distinction can be observed, between ‘inharmonic sounds which have little if any relevance for music (e.g., white or pink noise)’ (Schneider and Frieler,, 2009) and ‘coherent’ inharmonic signals, which ‘sound as stable and smooth as harmonic signals’ (de Boer,, 1956). In other words, inharmonicity can be the result of either inharmonic partial relations or noise. Section 6.6 elaborates on the relationship between noise and inharmonicity.
HR-inharmonicity and peak prominence are not independent. Adding white noise to a harmonic complex tone does not change the partial positions, yet it lowers the maximum value of the normalised auto-correlation after the first zero-crossing, resulting in lower HR-inharmonicity values (see section 6.6 for more details). To measure inharmonicity independently from ‘noisiness’, we perform a PCA on the distribution of HR-inharmonicity and peak prominence values that generates a 2-dimensional representation in which we denote the dimensions by PC1 and PC2. PC1 represents the total amount of noise and inharmonicity, while PC2 represents the proportion of this total that is due to inharmonic intervals between partials. In this paper, we refer to PC2 as inharmonicity.
@@ 3.2.4 Audio weighting
Figure 2 (a) shows the power spectrum values for the BEA dataset. These results are consistent with those of Pestana et al., (2013). However, these power spectrum values are not an accurate reflection of what the listener actually hears. Humans are not equally sensitive to all frequencies (Fletcher and Munson,, 1933). Several models exist for the level of tones that are perceived as equally loud by human listeners depending on their frequency and sound pressure level. One such model is ISO226:2003 (ISO,, 2003). Figure 2 (b) shows the power spectrum weighted using the ISO226:2003 equal-loudness contour corresponding to 50 phon, which is the median contour. Most notably, bass frequencies become attenuated. To more accurately reflect what the listener hears, we weigh the audio using the ISO226:2003 50 phon equal-loudness contour and evaluate the above features on both raw and weighted audio (see section 10 for more details).
4 Results
4.1 Measures of HR-inharmonicity
@@
Figure 3 shows HR-inharmonicity for the four datasets. As autocorrelation is robust to level, the tracks are first gated to remove low-level parts (extensive for instance in bonus tracks), which would result in very low feature values as a result of the analysis of the background noise. The gating process only retains the parts of the original audio whose RMS power is above -20dB after peak normalisation. From Figure 3, it is possible to make the following observations.
@@
In section 6.5, we point out that HR-inharmonicity is correlated with the music’s ‘roughness’ (in the sense of Masina et al.,, 2022).
4.2 Measures of peak prominence
] Refer to caption
@@ 4.2 Measures of peak prominence
Figure 4 shows peak prominence for the four datasets. From this graph, we can make the following observations.
@@
    Popular music and orchestral music are similarly noisy. From Figure 3, we derived that greater inharmonicity in popular music may stem from the use of inharmonic sources, more inharmonic musical intervals, and/or noise. We can now exclude noise from the causes. The greater HR-inharmonicity observed in popular music is therefore likely to derive from a greater use of complex tones with inharmonic partials as well as more frequent use of inharmonic intervals between complex tones.
@@
4.3 HR-inharmonicity and peak prominence: BEA dataset
Figure 5 (a) shows the result of the HR-inharmonicity vs. noisiness measures on the unweighted BEA dataset, including the evolution of both features according to the music’s year of release. For a single HR-inharmonicity value (same vertical position), it is possible to distinguish between inharmonicity that arises from noise (right-hand side of graph) or intervals between discrete components (left-hand side of graph). In Figure 5 (a), the bottom left corner of the graph corresponds to low noise and low HR-inharmonicity; whilst the top right corner corresponds to high noise and high HR-inharmonicity. The diagonal line extending from the bottom left of the graph to the top right corresponds to an increasing sum of noise and HR-inharmonicity. It is the axis that explains the most variance in the distribution.
Figure 5 (b) shows the result of the distribution’s PCA. In the upper two quadrants of the PCA graph, a relatively high proportion of the total amount of noise and inharmonicity is due to inharmonic intervals; while in the lower two quadrants, we have a relatively high proportion of the total amount of noise and inharmonicity being due to noise. Music tracks in the upper-right part of the PC1 vs. PC2 representation feature the highest amount of inharmonic relations between partials (high total amount of noise and inharmonicity, low proportion of this being due to noise). Conversely, tracks in the lower-left part feature the lowest density of inharmonic relations between partials.
@@ 4.3 HR-inharmonicity and peak prominence: BEA dataset
    1972–1986    During this period, there is generally increasing noise and increasing HR-inharmonicity with a relatively slightly greater increase in HR-inharmonicity, indicating that a slightly greater proportion of the HR-inharmonicity is due to inharmonic intervals, not noise, towards the end of this period.
@@
Figure 6: Respective positions of the four datasets. (a) HR-inharmonicity and noisiness. (b) PC1 and PC2 (as in Figure 5). The white, red, blue, and yellow stars respectively correspond to the BEA, orchestra, piano, and musique concrète datasets. The stars’ lines start at the median coordinates for each dataset, and stop where half of the points starting from the median were met.
4.4 HR-inharmonicity and peak prominence: all datasets
@@
    As we will discuss in more detail below in section 5.3, Musique concrète and CPM use similar ‘resources of the performance’ (Parry,, 1911), that is to say, they are both typically created in music production studios using electronic technology. In McPherson and Tahıroğlu,’s (2020) terms, musique concrète explores a noisier ‘space for musical exploration’ with fewer inharmonic intervals between partials, and CPM explores a less noisy space with more inharmonic intervals. Both are equally distant from orchestral music.
@@
4.5 HR-inharmonicity and peak prominence, weighted audio
@@
    The increase in the ratio of interval inharmonicity to noisiness (PC2) from 1961 to 1972 in Figure 5 is accompanied by a noticeable increase in total inharmonicity and noise (PC1) in Figure 7.
@@
    Between 1986 and 2000, the fall in inharmonicity to noise ratio (PC2) seems less than it does when viewed in terms of unweighted audio. It suggests that the faster evolution witnessed from non-weighted audio derives from very low frequencies. In other words, the faster evolution originates from a change in the properties of the medium in addition to that of the content. The weighting function gives relatively much less weight to the lower frequencies, meaning that the decrease in lower frequency energy between 1986 and 2000 would have less of an effect on the weighted audio graph of PC2.

### v2

@@ 3.2.2 Peak prominence and noisiness
Features have been proposed to measure the ‘noisiness’ of the signal, such as spectral flatness (Peeters,, 2004) and AudioSpectrumFlatnessType (ISO,, 2001). However, these two features are sensitive to the overall envelope of the spectrum: music and pink noise result in comparable spectral flatness values, which defeats the purpose of measuring ‘noisiness’. We introduce a new metric derived from spectral flatness and AudioSpectrumFlatnessType that is robust to the overall spectral profile. We refer to this feature as peak prominence and to the inverse of peak prominence as noisiness. Appendix A.2 details the elaboration of the metric.
@@ 3.2.3 Inharmonicity
HR-inharmonicity and peak prominence are not independent. Adding white noise to a harmonic complex tone does not change the partial positions, yet it lowers the maximum value of the normalised auto-correlation after the first zero-crossing, resulting in lower HR-inharmonicity values (see section A.1.6 for more details). To measure inharmonicity independently from ‘noisiness’, we perform a PCA on the distribution of HR-inharmonicity and peak prominence values that generates a 2-dimensional representation in which we denote the dimensions by PC1 and PC2. PC1 represents the total amount of noise and inharmonicity, while PC2 represents the proportion of this total that is due to inharmonic intervals between partials. In this paper, we refer to PC2 as inharmonicity.
@@ 3.2.4 Audio weighting
Figure 2 (a) shows the power spectrum values for the BEA dataset. These results are consistent with those of Pestana et al., (2013). However, these power spectrum values are not an accurate reflection of what the listener actually hears. Humans are not equally sensitive to all frequencies (Fletcher and Munson,, 1933). Several models exist for the level of tones that are perceived as equally loud by human listeners depending on their frequency and sound pressure level. One such model is ISO226:2003 (ISO,, 2003). Figure 2 (b) shows the power spectrum weighted using the ISO226:2003 equal-loudness contour corresponding to 50 phon, which is the median contour. Most notably, bass frequencies become attenuated. To more accurately reflect what the listener hears, we weigh the audio using the ISO226:2003 50 phon equal-loudness contour and evaluate the above features on both raw and weighted audio (see Appendix B for more details).
3.3 Corpus study
3.3.1 Measures of HR-inharmonicity
@@
Figure 3 shows HR-inharmonicity for the four datasets. As autocorrelation is robust to level, the tracks are first gated to remove low-level parts (e.g., extensively in bonus tracks), which would result in very low feature values as a result of the analysis of the background noise. The gating process only retains the parts of the original audio whose RMS power is above -20dB after peak normalisation. From Figure 3, it is possible to make the following observations.
@@
In Appendix A.1.5, we point out that HR-inharmonicity is correlated with the music’s ‘roughness’ (in the sense of Rasch and Plomp, (1982) and Masina et al., (2022)).
3.3.2 Measures of peak prominence
@@
Figure 4 shows the peak prominence for the four datasets. From this graph, we can make the following observations.
@@
    Popular music and orchestral music are similarly noisy. From Figure 3, we derived that greater inharmonicity in popular music may stem from the use of inharmonic sources, more inharmonic musical intervals, and/or noise. We can now exclude noise from the causes. The greater HR-inharmonicity observed in popular music is, therefore, likely to derive from a greater use of complex tones with inharmonic partials as well as the more frequent use of inharmonic intervals between complex tones.
@@
3.3.3 HR-inharmonicity and peak prominence: BEA dataset
Figure 5, left, shows the result of the HR-inharmonicity vs. noisiness measures on the unweighted BEA dataset, including the evolution of both features according to the music’s year of release. For a single HR-inharmonicity value (same vertical position), it is possible to distinguish between inharmonicity that arises from noise (right-hand side of graph) or intervals between discrete components (left-hand side of graph). In Figure 5, left, the bottom left corner of the graph corresponds to low noise and low HR-inharmonicity; whilst the top right corner corresponds to high noise and high HR-inharmonicity. The diagonal line extending from the bottom left of the graph to the top right corresponds to an increasing sum of noise and HR-inharmonicity. It is the axis that explains the most variance in the distribution.
Figure 5, right, shows the result of the distribution’s PCA. In the upper two quadrants of the PCA graph, a relatively high proportion of the total amount of noise and inharmonicity is due to inharmonic intervals; while in the lower two quadrants, we have a relatively high proportion of the total amount of noise and inharmonicity being due to noise. Music tracks in the upper-right part of the PC2-vs.-PC1 representation feature the highest amount of inharmonic relations between partials (high total amount of noise and inharmonicity, low proportion of this being due to noise). Conversely, tracks in the lower-left part feature the lowest density of inharmonic relations between partials.
@@ 3.3.3 HR-inharmonicity and peak prominence: BEA dataset
    1972–1986    During this period, there is generally increasing noise and increasing HR-inharmonicity with a relatively slightly greater increase in HR-inharmonicity, indicating that a slightly greater proportion of the increasing HR-inharmonicity is due to inharmonic intervals, not noise.
@@
Figure 6: Respective positions of the four datasets. (a) HR-inharmonicity and noisiness. (b) PC1 and PC2 (as in Figure 5). The white, red, blue, and yellow stars respectively correspond to the BEA, orchestra, piano, and musique concrète datasets. The stars’ lines start at the median coordinates for each dataset, and stop where half of the points starting from the median were met. The aquamarine-coloured curves at the centres of the images replicate the red curves in Figure 5.
3.3.4 HR-inharmonicity and peak prominence: all datasets
@@
    As we discussed in Section 2.2.3, Musique concrète and CPM use similar ‘resources of the performance’ (Parry,, 1911), that is to say, they are both typically created in music production studios using electronic technology. In McPherson and Tahıroğlu,’s (2020) terms, musique concrète explores a noisier ‘space for musical exploration’ with fewer inharmonic intervals between partials, and CPM explores a less noisy space with more inharmonic intervals. Both are equally distant from orchestral music.
@@
3.3.5 HR-inharmonicity and peak prominence, weighted audio
@@
    The increase in the ratio of interval inharmonicity to noisiness (PC2) from 1961 to 1972 in Figure 5 is accompanied by a noticeable increase in total inharmonicity and noise (PC1) in Figure 7 (b).
@@
    Between 1986 and 2000, the fall in inharmonicity to noise ratio (PC2) seems less than it does when viewed in terms of unweighted audio. It suggests that the faster evolution witnessed from non-weighted audio derives from very low frequencies. In other words, the faster evolution originates from a change in the properties of the medium in addition to that of the audible content. The weighting function gives relatively much less weight to the lower frequencies, meaning that the decrease in lower frequency energy between 1986 and 2000 would have less of an effect on the weighted audio graph of PC2 against PC1 (Figure 7 (b)).